Grade 12th passMechanics

A body is released from a height of 300m from the ground.Exactly at the same time another body is projected from the ground in verticaly upward direction with a velocity of 159m/s along the same line.Find when and where they will meet?

V G Munj
8 Years agoGrade 12th pass

1 Answer

Gaurav Gupta
8 Years ago
we can use the relative velocity approach here. let the body at a height of 300m is body A and body at ground is body B.
lets find velocity of A with respect to B.
VAB = VA-VB = 159 - 0 = 159m/s
similarly acceleration of A with respect to B 
aAB = aA-aB = -9.81-(-9.81) = 0
so now as B is at rest in this frame, A has to travel a distance of 300m to meet B, with “0” acceleration
so time taken to meet B will be        
 T=\frac{d}{V} = \frac{300}{159} = 1.88sec
so they will meet after 1.88sec
now coming back in ground frame distance travelled by block B in 1.88 sec
-S = uT-\frac{1}{2}gT^{2}  downwards is taken as negative
u =0
solving this gives
S = 17.5 m
so they will meet when B travells 17.5m or A travels 282.5m